OpenStudy (anonymous):

let g(x)= (x^2+x-6)/abs(x-2) ... find lim g(x) as x--> 2+ (from the right) and as x--> 2- (from the left)

OpenStudy (anonymous):

when we approach from right that is as x--->2+ abs(x-2) = (x-2) and when we approach from left that is as x--->2- abs(x-2)=-(x-2)

OpenStudy (anonymous):

that should help

OpenStudy (anonymous):

also x^2+x-6 = x^2+3x-2x-6 = x(x+3)-2(x+3)=(x-2)(x+3)

OpenStudy (anonymous):

thank you! I got it! :)